# Tackling Long-Horizon Tasks with Model-based Offline Reinforcement Learning

### v1

@@ Abstract
Model-based offline reinforcement learning (RL) is a compelling approach that addresses the challenge of learning from limited, static data by generating imaginary trajectories using learned models. However, it falls short in solving long-horizon tasks due to high bias in value estimation from model rollouts. In this paper, we introduce a novel model-based offline RL method, Lower Expectile Q-learning (LEQ), which enhances long-horizon task performance by mitigating the high bias in model-based value estimation via expectile regression of $\lambda$-returns. Our empirical results show that LEQ significantly outperforms previous model-based offline RL methods on long-horizon tasks, such as the D4RL AntMaze tasks, matching or surpassing the performance of model-free approaches. Our experiments demonstrate that expectile regression, $\lambda$-returns, and critic training on offline data are all crucial for addressing long-horizon tasks. Additionally, LEQ achieves performance comparable to the state-of-the-art model-based and model-free offline RL methods on the NeoRL benchmark and the D4RL MuJoCo Gym tasks.
@@ 1 Introduction
One of the major challenges in offline reinforcement learning (RL) is the overestimation of values for out-of-distribution actions due to the lack of environment interactions [21, 19]. Model-based offline RL addresses this issue by generating additional (imaginary) training data using a learned model, thereby augmenting the given offline data with synthetic experiences that cover out-of-distribution states and actions [34, 16, 35, 2, 30]. While these approaches have demonstrated strong performance in simple, short-horizon tasks, they struggle with noisy model predictions and value estimations, particularly in long-horizon tasks [23]. This challenge is evident in their poor performances (i.e. near zero) on the D4RL AntMaze tasks [6, 15].
@@
Typical model-based offline RL methods alleviate the inaccurate value estimation problem (mostly overestimation) by penalizing Q-values estimated from model rollouts with uncertainties in model predictions [34, 16] or value predictions [30, 14]. While these penalization terms prevent a policy from exploiting erroneous value estimations, the policy now does not maximize the true value, but maximizes the value penalized by heuristically estimated uncertainties, which can lead to sub-optimal behaviors. This is especially problematic in long-horizon, sparse-reward tasks, where Q-values are similar across nearby states [23].
@@
Another way to reduce bias in value estimates is by using multi-step returns [31, 12]. CBOP [14] constructs an explicit distribution of multi-step Q-values from thousands of model rollouts and uses this value as a target for training the Q-function. However, CBOP is computationally expensive for estimating a target value and uses multi-step returns solely for Q-learning, which provides insufficient learning signals for obtaining long-horizon behaviors.
@@
To tackle long-horizon tasks with model-based offline RL, we introduce a simple yet effective model-based offline RL algorithm, Lower Expectile Q-learning (LEQ). As illustrated in Figure 1, LEQ uses expectile regression with a small $\tau$ for both policy and Q-function training, providing an efficient and elegant way to achieve conservative Q-value estimates. Moreover, to better handle long-horizon tasks, we propose to optimize a policy and Q-function using $\lambda$-returns (i.e. TD($\lambda$) targets) of long ($15$-step) model rollouts, allowing the policy to directly learn from low-bias multi-step returns [28].
The experiments on the D4RL AntMaze and MuJoCo Gym tasks [6], as well as the NeoRL benchmark [26], demonstrate that our proposed conservative policy optimization with $\lambda$-return and critic training on offline data significantly improves offline RL policies in long-horizon tasks while achieving comparable performance in short-horizon, dense-reward tasks. Specifically, to the best of our knowledge, LEQ is the first model-based offline RL algorithm capable of matching or outperforming the performance of model-free offline RL algorithms on the long-horizon AntMaze tasks [6, 15].
@@
Figure 1: Lower Expectile Q-learning (LEQ). (left) In offline model-based RL, an agent can generate imaginary trajectories using a world model. (right) For conservative Q-evaluation of the policy, LEQ learns the lower expectile of the target $Q$-distribution from a sampled individual rollout $\mathcal{T}_{i}$, without estimating the entire Q-distribution with exhaustive rollouts.
@@ 2 Related Work
Offline RL [21] aims to solve a reinforcement learning problem only with pre-collected datasets, better than behavioral cloning policies [25]. One can simply apply off-policy RL algorithms on top of the fixed dataset. However, off-policy RL methods suffer from the overestimation of Q-values for actions unseen in the offline dataset [8, 18, 19], since an overestimated value function cannot get corrected through online environment interactions in offline RL.
@@
Model-free offline RL algorithms have addressed this value overestimation problem on out-of-distribution actions by (1) regularizing a policy to only output actions in the offline data [24, 17, 7] or (2) adopting a conservative value estimation for executing actions different from the dataset [19, 1]. Despite their strong performances on the standard offline RL benchmarks, model-free offline RL policies tend to be constrained to the support of the data (i.e. state-action pairs in the offline dataset), which may lead to limited generalization capability.
@@
Model-based offline RL approaches have tried to overcome this limitation by suggesting a better use of the limited offline data – learning a world model and generating imaginary data with the learned model that covers out-of-distribution actions. Similar to Dyna-style online model-based RL [32, 9, 10, 11], an offline model-based RL policy can be trained on both offline data and model rollouts. But, again, learned models may be inaccurate on states and actions outside the data support, making a policy easily exploit the learned models.
@@
Recent model-based offline RL algorithms have adopted the conservatism idea from model-free offline RL, penalizing policies incurring (1) uncertain transition dynamics [34, 16, 35] or (2) uncertain value estimation [30, 14]. This conservative use of model-generated data enables model-based offline RL to outperform model-free offline RL in widely used offline RL benchmarks [30]. However, uncertainty estimation is difficult and often inaccurate [35]. Instead of relying on such heuristic [34, 16, 30] or expensive [14] uncertainty estimation, we propose to learn a conservative value function via expectile regression with a small $\tau$, which is simple, efficient, yet effective.
@@ 3 Preliminaries
In this paper, we consider the offline RL setup [21], where a policy $\pi$ is trained with a fixed given offline dataset, $\mathcal{D}_{\text{env}}=\{\tau_{1},\tau_{2},...,\tau_{N}\}$, without any additional online interactions.
@@
As an offline RL policy is trained from a fixed dataset, one of the major challenges in offline RL is the limited data support; thus, lack of generalization to out-of-distribution states and actions. Model-based offline RL [16, 34, 35, 27, 30, 14] tackles this problem by augmenting the training data with imaginary training data (i.e. model rollouts) generated from the learned transition dynamics and reward model, $p_{\psi}(\mathbf{s}_{t+1},r\mid\mathbf{s}_{t},\mathbf{a}_{t})$.
The typical process of model-based offline RL is as follows: (1) pretrain a model (or an ensemble of models) and an initial policy from the offline data, (2) generate short imaginary rollouts $\{\tau\}$ using the pretrained model and add them to the training dataset $\mathcal{D}_{\text{model}}\leftarrow\mathcal{D}_{\text{model}}\cup\{\tau\}$, (3) perform an offline RL algorithm on the augmented dataset $\mathcal{D}_{\text{model}}\cup\mathcal{D}_{\text{env}}$, and repeat (2) and (3).
@@
Expectile is a generalization of the expectation of a distribution $X$. While the expectation of $X$, $\mathbb{E}[X]$, can be viewed as a minimizer of the least-square objective, $L_{2}(y)=\mathbb{E}_{x\sim X}[(y-x)^{2}]$, $\tau$-expectile of $X$, $\mathbb{E}^{\tau}[X]$, can be defined as a minimizer of the asymmetric least-square objective:
$L_{2}^{\tau}(y)=\mathbb{E}_{x\sim X}\left[\lvert\tau-\mathbbm{1}(y>x)\rvert\cdot(y-x)^{2}\right],$ (1)
where $\lvert\tau-\mathbbm{1}(y>x)\rvert$ is an asymmetric weighting of $L_{2}$ and $0\leq\tau\leq 1$.
We refer a $\tau$-expectile with $\tau<0.5$ as a lower expectile of $X$. When $\tau<0.5$, the objective assigns a high weight $1-\tau$ for smaller $x$ and a low weight $\tau$ for bigger $x$. Thus, minimizing the objective with $\tau<0.5$ leads to a conservative statistical estimate compared to the expectation.
@@ 4 Approach
The primary limitation for model-based offline RL in solving long-horizon tasks is inherent errors in a world model and critic outside the offline data support. Conservative value estimation can effectively handle such (falsely optimistic) errors. Prior approaches estimate conservative values through diverse uncertainty penalties; but they are either unreliable [35] or computationally expensive [14].
In this paper, we introduce Lower Expectile Q-learning (LEQ), an efficient model-based offline RL method that achieves conservative value estimation via expectile regression of Q-values with lower expectiles when learning from model-generated data (Section 4.1). Additionally, we address the noisy value estimation problem in long-horizon tasks [23] using $\lambda$-returns on $10$-step imaginary rollouts (Section 4.2). Finally, we train a deterministic policy conservatively by maximizing the lower expectile of $\lambda$-returns (Section 4.3). The overview of LEQ is described in Algorithm 1.
@@
3:  Pretrain $\pi_{\theta}$ and $Q_{\phi}$ on $\mathcal{D}_{\text{env}}$triangleright using BC for $\pi_{\theta}$ and FQE [20] for $Q_{\phi}$
@@
11:        $\mathbf{s}_{t+1},r_{t}\sim p_{\psi}(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$, where $p_{\psi}\sim\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$
12:     // Generate imaginary data, $\tau=\{(\mathbf{s}_{0},\mathbf{a}_{0},r_{0},\cdots,\mathbf{s}_{H-1},\mathbf{a}_{H-1},r_{H-1},\mathbf{s}_{H})_{i}\}$
@@
16:        $\mathbf{s}_{t+1},r_{t}\sim p_{\psi}(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$, where $p_{\psi}\sim\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$
17:     // Update critic using both offline and model-generated data
18:     Update critic $Q_{\phi}$ to minimize $\mathcal{L}^{\lambda}_{Q}(\phi)$ in Eq. 7 using $\tau$ and $\{\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime}\}\sim\mathcal{D}_{\text{env}}$
@@
20:     Update actor $\pi_{\theta}$ to minimize ^$\hat{\mathcal{L}}^{\lambda}_{\pi}(\theta)$ in Eq. 12 using $\tau$
@@ 4.1 Lower expectile Q-learning
As illustrated in Figure 1, the target value for $Q_{\phi}(\mathbf{s},\mathbf{a})$, where $\mathbf{a}\leftarrow\pi_{\theta}(\mathbf{s})$, can be estimated by rolling out an ensemble of world models and averaging $r(\mathbf{s},\mathbf{a})+\gamma Q_{\phi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})$ over all possible $\mathbf{s}^{\prime}$:
@@
This target value has three error sources: the predicted future state and reward $\mathbf{s}^{\prime},r\sim p_{\psi}(\cdot\mid\mathbf{s},\mathbf{a})$ and future Q-value $Q_{\phi}(\mathbf{s}^{\prime},\pi_{\theta}(\mathbf{s}^{\prime}))$. Thus, the target value computed from model-generate data, $\hat{y}_{\text{model}}$, is more prone to overestimation than that of the original target Q-value, $\hat{y}_{\text{env}}$, computed from $(\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime})\sim D_{\text{env}}$:
@@
To mitigate the overestimation problem in estimating the true Q-value from $H$-step inaccurate world model rollouts, we propose to use expectile regression on target Q-value estimation with small $\tau$; as illustrated in Figure 1, expectile regression with small $\tau$ tends to choose the target Q-value that is lower than the expectation, effectively providing a conservative estimate of target Q-value. Another advantage of using expectile regression is that we do not have to exhaustively evaluate Q-values to get $\tau$-expectiles as Jeong et al. [14]; instead, we can do conservative estimation using sampling:
$L_{Q,\text{model}}(\phi)=\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\tau\sim p_{\psi},\pi_{\theta}}\left[\frac{1}{H}\sum_{t=0}^{H}L_{2}^{\tau}(Q_{\phi}(\mathbf{s}_{t},\pi_{\theta}(\mathbf{s}_{t}))-\hat{y}_{\text{model}})\right].$ (4)
In addition, the Q-function is also trained on the offline data $\mathcal{D}_{\text{env}}$ with the standard Bellman update:
@@
To stabilize training of the Q-function, we adopt EMA regularization [11], which prevents drastic change of Q-values by regularizing the difference between the Q-predictions and ones from the exponential moving average:
$\mathcal{L}_{Q,\text{EMA}}(\phi)=\mathbb{E}_{(\mathbf{s},\mathbf{a})\in\mathcal{D}_{\text{env}}}\left[(Q_{\phi}(\mathbf{s},\mathbf{a})-Q_{\bar{\phi}}(\mathbf{s},\mathbf{a}))^{2}\right],$ (6)
where $\bar{\phi}$ is an exponential moving average of $\phi$. Note that by using EMA regularization, we do not use the target Q-network for Equations 2 and 3.
Finally, by combining the three losses above, we define the critic loss as follows:
$\mathcal{L}_{Q}(\phi)=\beta\mathcal{L}_{Q,\text{model}}(\phi)+(1-\beta)\mathcal{L}_{Q,\text{env}}(\phi)+\beta_{\text{EMA}}\mathcal{L}_{Q,\text{EMA}}(\phi).$ (7)

### v2

# Model-based Offline Reinforcement Learning with Lower Expectile Q-Learning

@@ Abstract
Model-based offline reinforcement learning (RL) is a compelling approach that addresses the challenge of learning from limited, static data by generating imaginary trajectories using learned models. However, these approaches often struggle with inaccurate value estimation from model rollouts. In this paper, we introduce a novel model-based offline RL method, Lower Expectile Q-learning (LEQ), which provides a low-bias model-based value estimation via lower expectile regression of $\lambda$-returns. Our empirical results show that LEQ significantly outperforms previous model-based offline RL methods on long-horizon tasks, such as the D4RL AntMaze tasks, matching or surpassing the performance of model-free approaches and sequence modeling approaches. Furthermore, LEQ matches the performance of state-of-the-art model-based and model-free methods in dense-reward environments across both state-based tasks (NeoRL and D4RL) and pixel-based tasks (V-D4RL), showing that LEQ works robustly across diverse domains. Our ablation studies demonstrate that lower expectile regression, $\lambda$-returns, and critic training on offline data are all crucial for LEQ.
@@ 1 Introduction
One of the major challenges in offline reinforcement learning (RL) is the overestimation of values for out-of-distribution actions due to the lack of environment interactions (Levine et al., 2020; Kumar et al., 2020). Model-based offline RL addresses this issue by generating additional (imaginary) training data using a learned model, thereby augmenting the given offline data with synthetic experiences that cover out-of-distribution states and actions (Yu et al., 2020; Kidambi et al., 2020; Yu et al., 2021; Argenson and Dulac-Arnold, 2021; Sun et al., 2023). While these approaches have demonstrated strong performance in simple, short-horizon tasks, they struggle with noisy model predictions and value estimates, particularly in long-horizon tasks (Park et al., 2024). This challenge is evident in their poor performances (i.e. near zero) on the D4RL AntMaze tasks (Fu et al., 2020).
@@
Typical model-based offline RL methods alleviate the inaccurate value estimation problem (mostly overestimation) by penalizing Q-values estimated from model rollouts with uncertainties in model predictions (Yu et al., 2020; Kidambi et al., 2020) or value predictions (Sun et al., 2023; Jeong et al., 2023). While these penalization terms prevent a policy from exploiting erroneous value estimates, the policy now does not maximize the true value, but maximizes the value penalized by heuristically estimated uncertainties, which can lead to sub-optimal behaviors. This is especially problematic in long-horizon, sparse-reward tasks, where Q-values are similar across nearby states (Park et al., 2024).
@@
Another way to reduce bias in value estimates is using multi-step returns (Sutton, 1988; Hessel et al., 2018). CBOP (Jeong et al., 2023) constructs an explicit distribution of multi-step Q-values from thousands of model rollouts and uses this value as a target for training the Q-function. However, CBOP is computationally expensive for estimating a target value and uses multi-step returns solely for Q-learning, not for policy optimization.
@@
To tackle these issues in model-based offline RL, we introduce a simple yet effective model-based offline RL algorithm, Lower Expectile Q-learning (LEQ). As illustrated in Figure 1, LEQ uses expectile regression with a small $\tau$ for both policy and Q-function training, providing an efficient and elegant way to achieve conservative Q-value estimates. Moreover, we propose to optimize both policy and Q-function using $\lambda$-returns (i.e. TD($\lambda$) targets) of long ($10$-step) model rollouts, allowing the policy to directly learn from low-bias multi-step returns (Schulman et al., 2016).
The experiments on the D4RL AntMaze, MuJoCo Gym (Fu et al., 2020), NeoRL (Qin et al., 2022), and V-D4RL (Lu et al., 2023) benchmarks show that LEQ improves model-based offline RL across diverse domains. To the best of our knowledge, LEQ is the first model-based offline RL algorithm capable of outperforming model-free offline RL algorithms on the long-horizon AntMaze tasks (Fu et al., 2020; Jiang et al., 2023). Moreover, LEQ matches the top scores across various benchmarks, while prior methods demonstrate superior performances only for a specific domain.
@@
Figure 1: Lower Expectile Q-learning (LEQ). (left) In model-based offline RL, an agent can generate imaginary trajectories using a world model. (right) For conservative Q-evaluation of the policy, LEQ learns the lower expectile of the target $Q$-distribution from a few sampled rollouts $\mathcal{T}_{i}$, without estimating the entire Q-distribution with exhaustive rollouts.
@@ 2 Related Work
Offline RL (Levine et al., 2020) aims to solve an RL problem only with pre-collected datasets, outperforming behavioral cloning policies (Pomerleau, 1989). While it is possible to apply off-policy RL algorithms on fixed datasets, these algorithms suffer from the overestimation of Q-values for actions unseen in the offline dataset (Fujimoto et al., 2019; Kumar et al., 2019; 2020) since the overestimated values cannot be corrected through interactions with environments as in online RL.
@@
Model-free offline RL algorithms have addressed this value overestimation problem on out-of-distribution actions by (1) regularizing a policy to only output actions in the offline data (Peng et al., 2019; Kostrikov et al., 2022; Fujimoto and Gu, 2021) or (2) adopting a conservative value estimation for executing actions different from the dataset (Kumar et al., 2020; An et al., 2021). Despite their strong performances on the standard offline RL benchmarks, model-free offline RL policies tend to be constrained to the support of the data (i.e. state-action pairs in the offline dataset), which may lead to limited generalization capability.
@@
Model-based offline RL approaches have tried to overcome this limitation by suggesting a better use of the limited offline data – learning a world model and generating imaginary data with the learned model that covers out-of-distribution actions. Similar to Dyna-style online model-based RL (Sutton, 1991; Hafner et al., 2019; 2021; 2023), an offline model-based RL policy can be trained on both offline data and model rollouts. But, again, learned models may be inaccurate on states and actions outside the data support, making a policy easily exploit the learned models.
@@
Recent model-based offline RL algorithms have adopted the conservatism idea from model-free offline RL, penalizing policies incurring (1) uncertain transition dynamics (Yu et al., 2020; Kidambi et al., 2020; Yu et al., 2021) or (2) uncertain value estimation (Sun et al., 2023; Jeong et al., 2023). This conservative use of model-generated data enables model-based offline RL to outperform model-free offline RL in widely used offline RL benchmarks (Sun et al., 2023). However, uncertainty estimation is difficult and often inaccurate (Yu et al., 2021). Instead of relying on such heuristic (Yu et al., 2020; Kidambi et al., 2020; Sun et al., 2023) or expensive (Jeong et al., 2023) uncertainty estimation, we propose to learn a conservative value function via expectile regression with a small $\tau$, which is simple, efficient, yet effective, as illustrated in Figure 2.
Expectile regression for offline RL has been first introduced by IQL (Kostrikov et al., 2022), which has been extended to model-based offline RL, such as IQL-TD-MPC (Chitnis et al., 2024). IQL uses “upper expectile” to approximate the “max operation” in $V(s)=\max_{a}Q(s,a)$ without querying out-of-distribution actions. On the other hand, our work fundamentally differs from IQL-like approaches in that our method uses “lower expectile” to get “conservative return estimates” from trajectories generated by potentially inaccurate model rollouts.
Q-learning Lower Expectile Q-learning
$\mathcal{L}_{Q}(\phi)=(Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t})-Q_{t}^{\lambda}(\mathcal{T}))^{2}$Longrightarrow ${\color[rgb]{0.1875,0.39453125,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.1875,0.39453125,0.6875}\lvert\tau-\mathbbm{1}(Q_{t}^{\lambda}(\mathcal{T})>Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t}))\rvert}\cdot\mathcal{L}_{Q}(\phi)$
$\nabla_{\theta}\mathcal{L}_{\pi}(\theta)=-\nabla_{\theta}\pi_{\theta}(\mathbf{s}_{t})\cdot\nabla_{\mathbf{a}_{t}}Q_{t}^{\lambda}(\mathcal{T})$Longrightarrow ${\color[rgb]{0.1875,0.39453125,0.6875}\definecolor[named]{pgfstrokecolor}{rgb}{0.1875,0.39453125,0.6875}\lvert\tau-\mathbbm{1}(Q_{t}^{\lambda}(\mathcal{T})>Q_{\phi}(\mathbf{s}_{t},\mathbf{a}_{t}))\rvert}\cdot\nabla_{\theta}\mathcal{L}_{\pi}(\theta)$
Figure 2: Comparison of standard Q-learning and Lower Expectile Q-learning (LEQ). LEQ generalizes standard Q-learning (with $\lambda$-returns $Q_{t}^{\lambda}(\mathcal{T})$) by multiplying a simple asymmetric weight $\lvert\tau-\mathbbm{1}(Q_{t}^{\lambda}(\mathcal{T})>Q_{\phi}(s_{t},a_{t}))\rvert$ to the Q-learning objectives. $\mathcal{T}=(\mathbf{s}_{0},\mathbf{a}_{0},r_{0},\mathbf{s}_{1},\mathbf{a}_{1},r_{1},\cdots,\mathbf{s}_{T})$ is a model-generated trajectory and $\tau\leq 0.5$ is the expectile hyperparameter controlling the degree of conservatism. When $\tau=0.5$, LEQ reduces to standard Q-learning.
@@ 3 Preliminaries
In this paper, we consider the offline RL setup (Levine et al., 2020), where a policy $\pi$ is trained with a fixed given offline dataset, $\mathcal{D}_{\text{env}}=\{\mathcal{T}_{1},\mathcal{T}_{2},...,\mathcal{T}_{N}\}$, without any additional online interactions.
@@
As an offline RL policy is trained from a fixed dataset, one of the major challenges in offline RL is the limited data support; thus, lack of generalization to out-of-distribution states and actions. Model-based offline RL (Kidambi et al., 2020; Yu et al., 2020; 2021; Rigter et al., 2022; Sun et al., 2023; Jeong et al., 2023) tackles this problem by augmenting the training data with imaginary training data (i.e. model rollouts) generated from the learned transition dynamics and reward model, $p_{\psi}(\mathbf{s}_{t+1},r_{t}\mid\mathbf{s}_{t},\mathbf{a}_{t})$.
The typical process of model-based offline RL is as follows: (1) pretrain a model (or an ensemble of models) and an initial policy from the offline data, (2) generate short imaginary rollouts $\{\mathcal{T}\}$ using the pretrained model and add them to the training dataset $\mathcal{D}_{\text{model}}\leftarrow\mathcal{D}_{\text{model}}\cup\{\mathcal{T}\}$, (3) perform an offline RL algorithm on the augmented dataset $\mathcal{D}_{\text{model}}\cup\mathcal{D}_{\text{env}}$, and repeat (2) and (3).
@@
Expectile is a generalization of the expectation of a distribution $X$. While the expectation of $X$, $\mathbb{E}[X]$, can be viewed as a minimizer of the least-square objective, $\mathbb{E}_{x\sim X}[L_{2}(y-x)]=\mathbb{E}_{x\sim X}[\frac{1}{2}(y-x)^{2}]$, $\tau$-expectile of $X$, $\mathbb{E}^{\tau}[X]$, can be defined as a minimizer of the asymmetric least-square objective $\mathbb{E}_{x\sim X}[L_{2}^{\tau}(y-x)]$, where $L_{2}^{\tau}(\cdot)$ is defined as:
$L_{2}^{\tau}(u)=\lvert\tau-\mathbbm{1}(u>0)\rvert\cdot u^{2},$ (1)
where $\lvert\tau-\mathbbm{1}(u>0)\rvert$ is an asymmetric weighting of least-squared objective with $0\leq\tau\leq 1$.
We refer to a $\tau$-expectile with $\tau<0.5$ as a lower expectile of $X$. When $\tau<0.5$, the objective assigns a high weight $1-\tau$ for smaller $x$ and a low weight $\tau$ for bigger $x$. Thus, minimizing the objective with $\tau<0.5$ leads to a conservative statistical estimate compared to the expectation.
@@ 4 Approach
The primary challenge of model-based offline RL is inherent errors in a world model and critic outside the support of offline data. Conservative value estimation can effectively handle such (falsely optimistic) errors. In this paper, we introduce Lower Expectile Q-learning (LEQ), an efficient model-based offline RL method that achieves conservative value estimation via expectile regression of Q-values with lower expectiles when learning from model-generated data (Section 4.1). Additionally, we address the noisy value estimation problem (Park et al., 2024) using $\lambda$-returns on $10$-step imaginary rollouts (Section 4.2). Finally, we train a deterministic policy conservatively by maximizing the lower expectile of $\lambda$-returns (Section 4.3). The overview of LEQ is described in Algorithm 1.
@@
3:  Pretrain $\pi_{\theta}$ and $Q_{\phi}$ on $\mathcal{D}_{\text{env}}$triangleright using BC for $\pi_{\theta}$ and FQE (Le et al., 2019) for $Q_{\phi}$
@@
11:        $\mathbf{s}_{t+1},r_{t}\sim p_{\psi}(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$, where $p_{\psi}\sim\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$triangleright sample $p_{\psi}$ every step
12:     // Generate imaginary data, $\mathcal{T}=\{(\mathbf{s}_{0},\mathbf{a}_{0},r_{0},\cdots,\mathbf{s}_{H-1},\mathbf{a}_{H-1},r_{H-1},\mathbf{s}_{H})_{i}\}$
@@
16:        $\mathbf{s}_{t+1},r_{t}\sim p_{\psi}(\cdot\mid\mathbf{s}_{t},\mathbf{a}_{t})$, where $p_{\psi}\sim\{p_{\psi_{1}},\cdots,p_{\psi_{M}}\}$triangleright sample $p_{\psi}$ every step
17:     // Update critic using both model-generated data and offline data
18:     Update critic $Q_{\phi}$ to minimize $\mathcal{L}^{\lambda}_{Q}(\phi)$ in Eq. 7 using $\mathcal{T}$ and $\{\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime}\}\sim\mathcal{D}_{\text{env}}$
@@
20:     Update actor $\pi_{\theta}$ to minimize ^$\hat{\mathcal{L}}^{\lambda}_{\pi}(\theta)$ in Eq. 11 using $\mathcal{T}$
@@ 4.1 Lower expectile Q-learning
As illustrated in Figure 1, the target value for $Q_{\phi}(\mathbf{s},\mathbf{a})$, where $\mathbf{a}\leftarrow\pi_{\theta}(\mathbf{s})$, can be estimated by rolling out an ensemble of world models and averaging $r+\gamma Q_{\phi}(\mathbf{s}^{\prime},\mathbf{a}^{\prime})$ over all possible $\mathbf{s}^{\prime}$:
@@
This target value has three error sources: the predicted future state and reward $\mathbf{s}^{\prime},r\sim p_{\psi}(\cdot\mid\mathbf{s},\mathbf{a})$ and future Q-value $Q_{\phi}(\mathbf{s}^{\prime},\pi_{\theta}(\mathbf{s}^{\prime}))$. Thus, the target value from model-generated data, $\hat{y}_{\text{model}}$, is more prone to overestimation than the original target Q-value, $\hat{y}_{\text{env}}$, computed from $(\mathbf{s},\mathbf{a},r,\mathbf{s}^{\prime})\sim D_{\text{env}}$:
@@
To mitigate the overestimation of $\hat{y}_{\text{model}}$ from inaccurate $H$-step model rollouts, we propose to use lower expectile regression on target Q-value estimates with small $\tau$. As illustrated in Figure 2, expectile regression with small $\tau$ learns a Q-function predicting Q-values lower than the expectation, i.e., a conservative estimate of a target Q-value. Another advantage of using lower expectile regression is that we do not have to exhaustively evaluate Q-values to get $\tau$-expectiles as Jeong et al. (2023); instead, we can learn a conservative Q-function with sampling:
$L_{Q,\text{model}}(\phi)=\mathbb{E}_{\mathbf{s}_{0}\in\mathcal{D}_{\text{model}},\mathcal{T}\sim p_{\psi},\pi_{\theta}}\left[\frac{1}{H}\sum_{t=0}^{H}L_{2}^{\tau}(Q_{\phi}(\mathbf{s}_{t},\pi_{\theta}(\mathbf{s}_{t}))-\hat{y}_{\text{model}})\right].$ (4)
Additionally, we train the Q-function with the transitions in the dataset $\mathcal{D}_{\text{env}}$, which do not have the risk of overestimation caused by the inaccurate model, using the standard Bellman update:
@@
To stabilize training of the Q-function, we adopt EMA regularization (Hafner et al., 2023), which prevents drastic change of Q-values by regularizing the difference between the predictions from $Q_{\phi}$ and ones from its exponential moving average $Q_{\bar{\phi}}$:
$\mathcal{L}_{Q,\text{EMA}}(\phi)=\mathbb{E}_{(\mathbf{s},\mathbf{a})\in\mathcal{D}_{\text{env}}}\left[(Q_{\phi}(\mathbf{s},\mathbf{a})-Q_{\bar{\phi}}(\mathbf{s},\mathbf{a}))^{2}\right].$ (6)
Finally, by combining the three aforementioned losses, we define the critic loss as follows:
$\mathcal{L}_{Q}(\phi)=\beta\mathcal{L}_{Q,\text{model}}(\phi)+(1-\beta)\mathcal{L}_{Q,\text{env}}(\phi)+\omega_{\text{EMA}}\mathcal{L}_{Q,\text{EMA}}(\phi).$ (7)